OpenStudy (anonymous):

Help! The unit for measuring work is Select one: a. J b. m c. N/m2 d. N

TheSmartOne (thesmartone):

Do you know the formula for work?

TheSmartOne (thesmartone):

@42cookiethere

TheSmartOne (thesmartone):

Okay then, \(\sf Work = Force \times Distance\) Do you know what the units for Force and the units for Distance would be?

OpenStudy (anonymous):

not really, sorry! haha! I really am not smart when it comes to this stuff!

OpenStudy (michele_laino):

hint: \[ \Large {\text{Work = Force }} \times {\text{ distance}}\]

TheSmartOne (thesmartone):

Well, look at the answer choices to figure out what the unit for force and the units for distance would be. Hint: N is newton m is meter

OpenStudy (anonymous):

@SkaterBoyShawn

OpenStudy (michele_laino):

hint: by definition, we have: \[\Large Joule{\text{ }} = 1Newton \cdot 1meter{\text{ }}\]